OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 9? 3 9 81

OpenStudy (anonymous):

what is the formula for a circle? x^2 + y^2 = r^2 ?

OpenStudy (anonymous):

its x^2 + y^2 =9

OpenStudy (anonymous):

radius is 3(3^2 = 9) just compare the equations

OpenStudy (anonymous):

yes, that's for this circle, but what about for any ole circle out there? Isn't it x^2 + y^2 = r^2 where r is the radius?

OpenStudy (anonymous):

for a circle centered at the origin, yeah

OpenStudy (anonymous):

so 9 would be the r^2; just solve that